The perimeter of parallelogram ABCD is 100 cm. AD is 5 cm more than twice AB. Find the lengths of all four sides of ABCD.
Answer by ewatrrr(24785)

 
Hi,

Parallelogram: Parallel Sides are Equal in Length
P = 2AD + 2AB = 100cm
Let x = AB, then AD = 2x + 5
Substitute:
2(2x+5) + 2x = 100cm
4x +10 + 2x = 100cm
6x = 90cm
x = 15cm and (2x+5) = 35cm
2*30cm + 2*35cm = 100cm Checks!
AB and DC have Lengths of 15cm
AD and BC have Lengths of 35cm
Wish You the Best in your Studies.